# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ 5 Properties of Least Squares Quasiconformal Energy
$\mathcal{M}_{f}^{*}\mathcal{M}_{p}\mathbf{U}_{p}$ $=\mathcal{N}_{f}^{*}\mathcal{N}_{p}\mathbf{U}_{p}+f_{1}\mathbf{v}_{1}\begin{pmatrix}\mathbf{0}\\ f_{2}^{*}\\ \mathbf{0}\\ f_{3}^{*}\\ \mathbf{0}\end{pmatrix},\mathcal{M}_{f}^{*}\mathcal{M}_{f}\mathbf{U}_{f}=\mathcal{N}_{f}^{*}\mathcal{N}_{f}\mathbf{U}_{f}+(f_{2}\mathbf{v}_{2}+f_{3}\mathbf{v}_{3})\begin{pmatrix}\mathbf{0}\\ f_{2}^{*}\\ \mathbf{0}\\ f_{3}^{*}\\ \mathbf{0}\end{pmatrix}$
@@
$\qquad(\sqrt{\alpha_{1}}\sum_{j}f_{j}\mathbf{v}_{j})l_{12}^{*}+(l_{32}\mathbf{v}_{2}+p_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j}+l_{31}\mathbf{v}_{1})l_{32}^{*}$
$=(\sum_{j}f_{j}\mathbf{v}_{j})(\alpha_{1}f_{2}^{*}-\alpha_{2}f_{1}^{*})+\frac{1}{\alpha_{3}}(\alpha_{3}f_{2}^{*}-\alpha_{2}f_{3}^{*})\cdot\left[(\alpha_{3}f_{2}-\alpha_{2}f_{3})\mathbf{v}_{2}+f_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j}+(\alpha_{3}f_{1}-\alpha_{1}f_{3})\mathbf{v}_{1}\right]$
$=(\sum_{j}f_{j}\mathbf{v}_{j})\left[(\alpha_{1}f_{2}^{*}-\alpha_{2}f_{1}^{*})+(\alpha_{3}f_{2}^{*}-\alpha_{2}f_{3}^{*})\right]=(\sum_{j}f_{j}\mathbf{v}_{j})(\alpha_{1}f_{2}^{*}+\alpha_{2}f_{2}^{*}+\alpha_{3}f_{2}^{*})=(\sum_{j}f_{j}\mathbf{v}_{j})f_{2}^{*}$
@@
$\qquad p_{1}^{*}(\sqrt{\alpha_{1}}\sum_{j}f_{j}\mathbf{v}_{j})+p_{2}^{*}(l_{23}\mathbf{v}_{3}+p_{2}\sum_{j}\alpha_{j}\mathbf{v}_{j})+p_{3}^{*}(l_{32}\mathbf{v}_{2}+p_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j})+l_{21}\mathbf{v}_{1}p_{2}^{*}+l_{31}\mathbf{v}_{1}p_{3}^{*}$

### v2

@@ 3 Properties of Least-squares Quasiconformal Energy
$\mathcal{M}_{f}^{*}\mathcal{M}_{p}\mathbf{U}_{p}$ $=\mathcal{N}_{f}^{*}\mathcal{N}_{p}\mathbf{U}_{p}+f_{1}\mathbf{v}_{1}\begin{pmatrix}\mathbf{0}\\ f_{2}^{*}\\ \mathbf{0}\\ f_{3}^{*}\\ \mathbf{0}\end{pmatrix}$
@@
$\mathcal{M}_{f}^{*}\mathcal{M}_{f}\mathbf{U}_{f}$ $=\mathcal{N}_{f}^{*}\mathcal{N}_{f}\mathbf{U}_{f}+(f_{2}\mathbf{v}_{2}+f_{3}\mathbf{v}_{3})\begin{pmatrix}\mathbf{0}\\ f_{2}^{*}\\ \mathbf{0}\\ f_{3}^{*}\\ \mathbf{0}\end{pmatrix}$
${\mathcal{M}^{+}_{f}}^{*}{\mathcal{M}^{+}_{f}}\begin{pmatrix}\mathbf{U}_{f}\\ \sum_{j}\alpha_{j}\mathbf{v}_{j}\end{pmatrix}$ $={\mathcal{M}_{f}^{+}}^{*}\begin{pmatrix}\mathcal{N}_{f}\mathbf{U}_{f}\\ l_{12}\mathbf{v}_{2}+l_{13}\mathbf{v}_{3}+p_{1}\sum_{j}\alpha_{j}\mathbf{v}_{j}\\ l_{23}\mathbf{v}_{3}+p_{2}\sum_{j}\alpha_{j}\mathbf{v}_{j}\\ l_{32}\mathbf{v}_{2}+p_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j}\end{pmatrix}$
@@
$\quad(\sqrt{\alpha_{1}}\sum_{j}f_{j}\mathbf{v}_{j})l_{12}^{*}+(l_{32}\mathbf{v}_{2}+p_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j}+l_{31}\mathbf{v}_{1})l_{32}^{*}$
$=(\sum_{j}f_{j}\mathbf{v}_{j})(\alpha_{1}f_{2}^{*}-\alpha_{2}f_{1}^{*})+\frac{1}{\alpha_{3}}(\alpha_{3}f_{2}^{*}-\alpha_{2}f_{3}^{*})\cdot$
$\quad\left[(\alpha_{3}f_{2}-\alpha_{2}f_{3})\mathbf{v}_{2}+f_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j}+(\alpha_{3}f_{1}-\alpha_{1}f_{3})\mathbf{v}_{1}\right]$
$=(\sum_{j}f_{j}\mathbf{v}_{j})\left[(\alpha_{1}f_{2}^{*}-\alpha_{2}f_{1}^{*})+(\alpha_{3}f_{2}^{*}-\alpha_{2}f_{3}^{*})\right]$
$=(\sum_{j}f_{j}\mathbf{v}_{j})(\alpha_{1}f_{2}^{*}+\alpha_{2}f_{2}^{*}+\alpha_{3}f_{2}^{*})=(\sum_{j}f_{j}\mathbf{v}_{j})f_{2}^{*}$
@@
$\quad p_{1}^{*}(\sqrt{\alpha_{1}}\sum_{j}f_{j}\mathbf{v}_{j})+p_{2}^{*}(l_{23}\mathbf{v}_{3}+p_{2}\sum_{j}\alpha_{j}\mathbf{v}_{j})+p_{3}^{*}(l_{32}\mathbf{v}_{2}+$
$\quad\quad p_{3}\sum_{j}\alpha_{j}\mathbf{v}_{j})+l_{21}\mathbf{v}_{1}p_{2}^{*}+l_{31}\mathbf{v}_{1}p_{3}^{*}$